What is the potential difference required to stop a proton traveling at an initial speed of 2.33 x 107 m/s? ( mass of a proton is 1.67 x 10-27 kg)
1
Expert's answer
2020-03-02T09:09:18-0500

The kinetic energy = "\dfrac{1}{2}mv^2" ="0.5 \times1.67 \times10^{-27} \times(2.33 \times10^7)^2" ="4.53 \times10^{-13} J"

charge* potential difference = kinetic energy

"1.6 \times10^{-19} \times V=4.53 \times10^{-13} \newline potential \ difference =2.83 \times10^{6} Volts"